OpenStudy (anonymous):

A plumber charges $55 per hour plus a $50 service charge. Kevin hired the plumber to install a new hot water heater. What are the total charges if the plumber took 4 hours to complete the installation?

OpenStudy (anonymous):

50 + h(55) = 50 + 4(55) = 270